# Representing Matroids over the Reals is $\exists\mathbb{R}$-complete

### v3

@@ 2 Proof Overview and Background
Therefore, $\mathbb{F}$-representability appears to be in NP $\cap$ co-NP when the input is given as the exhaustive list of bases for each finite $\mathbb{F}$ given the claimed proof of Rota’s conjecture. Given that, deciding the computational complexity of $\mathbb{F}$-representability for each finite $\mathbb{F}$ with explicit bases description is an intriguing question. For $\mathbb{F}=GF(2),$ a polynomial-time algorithm is straightforward from the uniqueness of a binary representation (up to linear transformation) and the fact that such a representation can be efficiently obtained [51]; after constructing a matrix over $GF(2)$, we apply the above verification algorithm for NP membership. However, even for $\mathbb{F}=GF(3)$ it is not clear whether a matrix over GF(3) can be efficiently constructed although it is known that there is a unique representation over $GF(3)$ for a matroid representable over $GF(3).$ As far as we are aware, there is no efficient procedure known for constructing the representation of a matroid $M$ with a promise that $M$ is representable over $GF(3),$ when $M$ is given with an independence oracle or even given as a matrix over the rationals $\mathbb{Q}$ [30]. Getting an input matroid as explicit bases description might help to circumvent this obstacle.
@@
In contrast to the case of finite fields, it is impossible to certify $-\mathbb{R}$-representability with a polynomial number of rank evaluations [51]. Finally, for $\mathbb{R}$-representability we showed $\exists\mathbb{R}$-completeness, exhibiting a noticeable diversion from $\mathbb{F}$-representability for finite $\mathbb{F}$ which appears neither NP-complete nor co-NP-complete with explicit bases description under the assumption NP $\neq$ co-NP. Therefore, our result highlights the recurring contrast between representability over a finite field and over the reals.
@@ 3 Distinct-ETR
Note that steps 1, 2, and 3 have already been done (among others) by Schaefer and Štefankovič [56]. We sketch the main steps of their reduction. Specifically, we point out some properties that were not explicitly emphasized. We start to explain a simple trick that is excessively used in those types of constructions in order to build small, very small, and very large numbers.
@@
We define the problem ETRAMI (see [4] for more background on this problem) as a variant of ETR as follows. We are given a set of variables $X=\{x_{1},\ldots,x_{n}\}$ and constraints of the form
@@
for $x,y,z\in X.$ Therefore, ETRAMI is a variant of ETR without negations, inequalities, disjunctions, conjunctions and the only constant is $1$. It is folklore that ETRAMI is $\exists\mathbb{R}$-complete and follows implicitly from various papers [45, 56, 60]
@@
Again, $\exists\mathbb{R}$-membership follows from the fact that Feasibility is a special case of ETR. To show hardness we sketch a reduction from ETRAMI that is already known [45, 56]. Let $f_{1}=0,\ldots,f_{m}=0$ be the equations of some ETRAMI instance $\varphi$. (For example $x+y=z$ becomes $x+y-z=0$.) Let
@@
Again membership follows from the fact that ETR is more general then STRICT-INEQ. To show $\exists\mathbb{R}$-hardness we reduce from Feasibility. The idea of the reduction is to replace $p(x)=0$ by $-\delta<p(x)<\delta$, for some sufficiently small $\delta$. To this end, we employ two lemmas as formulated by Schaefer and Štefankovič [56]. Note that the actual proof comes from real algebraic geometry and can be found for instance in [6] and [33].
@@ 4 Arithmetic Using Matroids
In this section, we describe how to encode addition and multiplication of real numbers using rank-$3$ matroids. We rely on the von Staudt constructions, which are very well-known, perhaps with the caveat that they are usually stated for oriented matroids. But as we shall see, no orientedness is actually required to make them work. We follow the presentation of Matoušek [45].
@@ 5 Proof of Theorem 2
We first consider the case of rank equal to $3$ and treat the general case later. Let $O=(E,\chi)$ be an order type on $n$ elements of rank $3$. (We assume that $E=\{e_{1},\ldots,e_{n}\}$.) We construct a matroid $M$ from $O$ inductively. To be more precise, we construct matroids, $M_{3},M_{4},\ldots,M_{n}=M$ such that $M_{i}$ simulates $O_{i}$. Here, $O_{i}=(E_{i},\chi_{i})$ is the order type formed by the first $i$ elements of $O$. All of our matroids $M_{i}$ will feature a distinguished line $\ell_{\infty}$.

### v4

@@ 2 Proof Overview and Background
Therefore, $\mathbb{F}$-representability appears to be in NP $\cap$ co-NP when the input is given as the exhaustive list of bases for each finite $\mathbb{F}$ given the claimed proof of Rota’s conjecture. Given that, deciding the computational complexity of $\mathbb{F}$-representability for each finite $\mathbb{F}$ with explicit bases description is an intriguing question. For $\mathbb{F}=GF(2),$ a polynomial-time algorithm is straightforward from the uniqueness of a binary representation (up to linear transformation) and the fact that such a representation can be efficiently obtained (Oxley, 2006); after constructing a matrix over $GF(2)$, we apply the above verification algorithm for NP membership. However, even for $\mathbb{F}=GF(3)$ it is not clear whether a matrix over GF(3) can be efficiently constructed although it is known that there is a unique representation over $GF(3)$ for a matroid representable over $GF(3).$ As far as we are aware, there is no efficient procedure known for constructing the representation of a matroid $M$ with a promise that $M$ is representable over $GF(3),$ when $M$ is given with an independence oracle or even given as a matrix over the rationals $\mathbb{Q}$ (Hlinený, 2006). Getting an input matroid as explicit bases description might help to circumvent this obstacle.
@@
In contrast to the case of finite fields, it is impossible to certify $-\mathbb{R}$-representability with a polynomial number of rank evaluations (Oxley, 2006). Finally, for $\mathbb{R}$-representability we showed $\exists\mathbb{R}$-completeness, exhibiting a noticeable diversion from $\mathbb{F}$-representability for finite $\mathbb{F}$ which appears neither NP-complete nor co-NP-complete with explicit bases description under the assumption NP $\neq$ co-NP. Therefore, our result highlights the recurring contrast between representability over a finite field and over the reals.
@@ 3 Distinct-ETR
Note that steps 1, 2, and 3 have already been done (among others) by Schaefer and Štefankovič (2017). We sketch the main steps of their reduction. Specifically, we point out some properties that were not explicitly emphasized. We start to explain a simple trick that is excessively used in those types of constructions in order to build small, very small, and very large numbers.
@@
We define the problem ETRAMI (see Abrahamsen and Miltzow (2019) for more background on this problem) as a variant of ETR as follows. We are given a set of variables $X=\{x_{1},\ldots,x_{n}\}$ and constraints of the form
@@
for $x,y,z\in X.$ Therefore, ETRAMI is a variant of ETR without negations, inequalities, disjunctions, conjunctions and the only constant is $1$. It is folklore that ETRAMI is $\exists\mathbb{R}$-complete and follows implicitly from various papers, for example Matoušek (2014); Schaefer and Štefankovič (2017); Shor (1991).
@@
Again, $\exists\mathbb{R}$-membership follows from the fact that Feasibility is a special case of ETR. To show hardness we sketch a reduction from ETRAMI that is already known (Matoušek, 2014; Schaefer and Štefankovič, 2017). Let $f_{1}=0,\ldots,f_{m}=0$ be the equations of some ETRAMI instance $\varphi$. (For example $x+y=z$ becomes $x+y-z=0$.) Let
@@
Again membership follows from the fact that ETR is more general then STRICT-INEQ. To show $\exists\mathbb{R}$-hardness we reduce from Feasibility. The idea of the reduction is to replace $p(x)=0$ by $-\delta<p(x)<\delta$, for some sufficiently small $\delta$. To this end, we employ two lemmas as formulated in Schaefer and Štefankovič (2017). Note that the actual proof comes from real algebraic geometry and can be found for instance in Basu and Roy (2010) and Jeronimo, Perrucci, and Tsigaridas (2013).
@@ 4 Arithmetic Using Matroids
In this section, we describe how to encode addition and multiplication of real numbers using rank-$3$ matroids. We rely on the von Staudt constructions, which are very well-known, perhaps with the caveat that they are usually stated for oriented matroids. But as we shall see, no orientedness is actually required to make them work. We follow the presentation of Matoušek (2014).
@@ 5 Proof of Theorem 2
We first consider the case of rank equal to $3$ and treat the general case later. Let $O=(E,\chi)$ be an order type on $n$ elements of rank $3$. (We assume that $E=\{e_{1},\ldots,e_{n}\}$.) We construct a matroid $M$ from $O$ inductively. To be more precise, we construct matroids, $M_{3},\ldots,M_{n}=M$ such that $M_{i}$ simulates $O_{i}$. Here, $O_{i}=(E_{i},\chi_{i})$ is the order type formed by the first $i$ elements of $O$. All of our matroids $M_{i}$ will feature a distinguished line $\ell_{\infty}$.
@@ 6 Proof of Theorem 1
Acknowledgements.
We are grateful to the anonymous reviewers for their numerous suggestions that greatly improved the article.